9. 2/10 points us Answers 1-D Kinematics: Free Fall- Word Prob (v & v o) My Notes Ask Your Te A child on a bridge throws a rock straight down to the water below. The point where the child released the rock is 91 m above the water and it took 0.8 s for the rock to reach the water. Determine the rocks velocity (magnitude & direction) at the moment the child released it. Also determine the rocks velocity (magnitude & direction) at the moment it reached the water. Ignore air drag. ที่.-12.01 796 m/s X |down +0m/s down 10. 10/10 points | Previous Answers 1-D Kinematics: Free Fall- Word Prob (y & t) A cannonball is fired straight up into the air. The cannonball leaves the cannon at a speed of 73 m/s. Determine how high the cannonball goes and how long it took to reach the peak. Ignore air drag. Max height 271.6 m Time 7.441s My Notes Ask Your T 11. 5/10 points 1 Previous Answers 1-D Kinematics: New Computational Problem (v1) My Notes Ask Your Teacher A UFO experienced a constant acceleration of 6 m/s2 north during which its displacement was 20,485 m north and was going 496 m/s north at the end of said acceleration. Determine how long the UFO was accelerating & the initial speed of the UFO. NOTE: Initially the UFO was going south. At 20244 s V14 m/s Submit Assignment

9 vi and vf and 11 delta t and please show formulas and round 4 significant figures. Thank you
Answer #1

Solution) This problem is based on kinematics .

In kinematics we mainly use three formula and these are the basic formulas

(1) V = U + at

(2) S = Ut + (1/2)at^2

(3) V^2 = U^2 + 2aS

Here V is final velocity otherwise represented as Vf and U is initial velocity otherwise represented as Vi

S is displacement

a is acceleration due to gravity otherwise termed as g

For a body falling down a= g = 9.8m/s^2

For a body moving up a = g = -9.8m/s^2

t is time

olun) m a Rnilads and Vi (down) This aline L。dy asit o released down The rocks vely ma at Momenナ it reached water, V we havt C3 Ve 42.23 on ms dowon mls down 스t at ะ 416-14 -80-373